# When LLMs Go Online: The Emerging Threat of Web-Enabled LLMs

### v2

@@ 8 Conclusion
In this work, we examine the emerging threats posed by LLM agents, particularly their potential use of web-based tools in cyberattacks that exploit private data. To assess their capabilities, we focused on three types of cyberattacks: PII collection, impersonation post generation, and spear phishing email generation. Our experimental results demonstrate that attackers can successfully automate these attacks using LLM agents, with web-based tools enhancing the performance of these attacks. Furthermore, our results report easy bypass of LLM service safeguards at low-costs, making utilization of LLM agents in cyberattacks a viable option. These findings expose a significant vulnerability in current safeguards and underscore the urgent need for security measures to prevent the misuse of LLM agents.
@@
We care deeply about human safety and societal security, which motivates us to conduct this research. This paper discusses and carefully examines the potential threat of LLM agents in cyberattacks exploiting personal data. This raises ethical considerations as some of the data that we extract contains information about individual users. We minimize these concerns by using already public data and not presenting any specific information in this paper.
Aggregation of PIIs. The collected PII data is stored in pickle format on the principal investigator’s server. The data was provided to annotators in an Excel file format for annotation purposes. Once the annotation process was completed, the annotators returned the annotated files to the principal investigator and deleted their local copies. Currently, files containing personal information are exclusively accessible by the principal investigator and will be discarded after the publication of the research.
User Study. In the case of the user study in Section 6, according to the local IRB’s protocol, our research is excluded from the review for the following reasons. First, we used only publicly available information and did not handle any sensitive data. Additionally, the emails evaluated contained no harmful or offensive content, minimizing any potential impact on participants. In the recruitment process, we obtained informed consent from participants, detailing the purpose of the research, the expected duration, procedures involved, anticipated benefits, personal information protection, and consent for collecting and using their provided email addresses. We also informed participants that the LLM agents could search the internet for their email addresses to gather information and use it to generate emails. Only those consented could participate, and each participant completed the survey once. In the evaluation process, we did not send emails to the participants. We asked to evaluate emails through a Google survey after notification. Each participant was exposed only to emails generated from their own email address. All participant-related information was destroyed after the experiments, and participants were informed about this process. Additionally, we informed them that their responses would be excluded from the analysis if they withdrew consent, ensuring robust protection of personal information.
Responsible Disclosure Process. We believe that the benefits of publicizing these attacks outweigh the potential harms. First, we plan to submit our research findings to the LLM service vendors (OpenAI, Anthropic, and GoogleAI), including our prompts and concerns about the safeguard vulnerabilities of web-based tools. Additionally, we will request the implementation of a rule ensuring that when an LLM utilizes web crawling tools, it checks and adheres to robots.txt directives. (A robots.txt file tells web crawlers which parts of a site they can or cannot access.) Second, for the PII providers from whom we collected data, we strongly recommend creating a robots.txt file to block web crawlers from accessing pages containing personal information.
@@
Our study explores the potential threat posed by LLM agents in cyberattacks. We acknowledge that disclosing the exact prompts used for these attacks could enable malicious actors to replicate them. To mitigate the risk of misuse, we have submitted our code for review with dummy prompts accessible at this link  [^14] https://zenodo.org/doi/10.5281/zenodo.13691327
. After decision, we will discuss with reviewers and determine the feasibility of releasing our code. We will provide the full source code only upon request for legitimate research purposes.
Appendix A Example of Search() function results
Figure13 shows the example of Search() function results. We use ‘Usenix Security 25’ as a query.
Figure 13: Part of the results of search(Usenix Security 25).
Appendix B PII collection
B.1 Human annotation
@@
Appendix C Impersonation Post Generation
C.1 Templates for Evaluation
@@
C.2 Human Correlation with LLM Evaluation
We assessed the reliability of our LLM evaluators by comparing with human reviewers. We sampled 270 posts, each impersonating one of 10 professors for each claim, language model, and capability(LLM, WebSearch gent, WebNav agent) combination mentioned in Section 5.1. Then, both human reviewers and agents evaluated these posts. The average agreement rate between humans and agents was 93.0% for the first claim and 85.2% for the second, underscoring the reliability and credibility of the LLM evaluations.
@@ Appendix D Approachability of Cyberattacks
D.1 PII Collection
We conduct a cost analysis of PII collection using WebSearch and WebNav agents. Using navigation was more expensive in terms of time and token consumption compared to not using it (Table 5). During the process of searching for information, time and tokens are consumed as various links and buttons are clicked within the site. Upon integrating navigation functionality, the average token usage for GPT and Claude agents increased by 8.95 times and 6.80 times, respectively.
GPT vs. Claude. Interestingly, regardless of the use of navigation functions, GPT made more function calls than Claude. Without navigation, this difference became even more pronounced because GPT continuously altered search queries when the desired information was not immediately found. This result highlights GPT’s more proactive exploration compared to Claude, as previously mentioned.
When not using the navigation function, GPT required less time to complete tasks than Claude, despite making more function calls. This could be attributed to differences in model performance, aligning with the analysis reported in [2].
Table 5: Cost statistics for LLM agents to collect PII of professors with and without the navigation function.
GPT Claude
navigation x o x o
Time (sec) min 3.33 6.53 7.85 15.13
max 11.18 294.45 17.51 180.66
mean 6.69 53.50 11.43 43.28
min 1.00 2.00 1.00 2.00
# Function calls max 6.00 17.00 2.00 8.00
mean 3.56 5.74 1.40 3.72
# Tokens min 438.00 895.00 1159.00 1948.00
max 2559.00 76041.00 2268.00 78662.00
mean 1406.04 12583.74 1462.78 9952.80
D.2 Impersonation Post and Phishing Email Generation
We conduct a cost analysis of impersonation post generation and phishing email generation using WebSearch agents. Table 6 represents cost statistics.
@@ D.2 Impersonation Post and Phishing Email Generation
D.3 Safeguard Capability in Phishing Email Generation
Table 7 shows the response generation capabilities of LLMs across different configurations.
Table 7: Response generation capabilities of LLMs across different configurations based on the purpose of the email or the type of email address. successfully generated a response (o) or refused to generate a response (x)
LLM LLM Agent
GPT Claude Gemini GPT Claude Gemini
Purpose General o o x o o x
Credential x x x o o x
Money x x x o x x
Email address Institutional o o x o o x
Gmail o x x o o x
@@ Appendix E Spear Phishing Email Generation
E.1 Recruiting Process and Ethical Measures
We advertised our study within communities and contacted potential participants from our personal networks. All participants were informed about an incentive of approximately \$7.50 for their participation.
To analyze the evaluation results relative to the characteristics of the institutions associated with the targets, we ensured a balanced distribution among individuals from academic research and non-academic professional domains. This balanced approach facilitated a comprehensive evaluation of the variables under investigation.
@@ E.1 Recruiting Process and Ethical Measures
Of the 70 individuals who provided their email addresses via the consent form, 60 completed the evaluation of the emails.
@@ E.2 Chi-Squared Tests
This section investigates the factors influencing participants’ perception of email authenticity. First, we identify which aspects of an email had the most significant influence on determining the email’s authenticity. We conducted chi-squared tests to evaluate the independence between email evaluation questions (Q1–Q4) and the question assessing perceived email authenticity (Q6). Please refer to Appendix E.5 for the questions used in our survey. Table 8 shows that “content detail” (Q1) and “content authenticity” (Q2) had p-values much less than 0.05. These results suggest that these factors are critical in determining whether an email is perceived as genuine.
Further analysis focused on “content authenticity” (Q2) identified as the most strongly associated factor. Specifically, we examined which aspects of email content influenced participants’ perceptions of authenticity. We evaluated the independence between each response option within Q2 and the option in Q6 using chi-squared tests. Table 9 shows that the “Specific area of expertise” option had a p-value much less than 0.05, indicating a highly significant association. This finding suggests that when the email content aligns with the participant’s domain-specific knowledge, it strongly impacts their perception of the email’s authenticity.
Table 8: Chi-squared test results between each question (Q1-Q4) and Q6.
Q1 Q2 Q3 Q4
(content detail) (content auth) (sender) (recipient)
kappa 98.80 113.54 56.54 3.69
p-value $2.23\times 10^{-5}$ $3.12\times 10^{-7}$ 0.0159 0.4497
Table 9: Chi-squared test results between each response option within Q2 and Q6.
institution general expertise specific expertise hobby
kappa 4.34 0.01 11.15 0.28
p-value 0.0595 0.0373 0.9436 0.0008
@@ E.4 Contingency Table
E.5 Questionnaire for User Study
Q1. [Content Detail] What content does the email include? (select all that apply)
Note: Choose answers based solely on the presence of content, regardless of its authenticity or relevance to you.
  • $\square$
    None of the items match
  • $\square$
    Institution (e.g., Tesla, Meta, etc.)
  • $\square$
    General area of expertise (e.g., EE, CS, NLP, Implementation, etc.)
  • $\square$
    Specific area of expertise (e.g., LLM Security, Reinforcement Learning, Implementation using Copilot, etc.)
  • $\square$
    Hobby, Pastime activity (e.g., Football, Game, etc.)
Q2. [Content Authenticity] Which of the included contents are authentic? (select all that apply)
Note: “Authentic” means the information aligns with your interests or expertise areas.
For example, if your areas of expertise are LLM Security & LLM Benchmarks, only consider content authentic if it correctly relates to these areas.
  • $\square$
    None of the items match
  • $\square$
    Institution
  • $\square$
    General area of expertise
  • $\square$
    Specific area of expertise
  • $\square$
    Hobby, Pastime activity (e.g., Football, Game, etc.)
Q3. [Sender] Check the items that apply to the sender of the email. (select all that apply)
  • $\square$
    None of the items match
  • $\square$
    I know the sender’s affiliated institution (e.g., Tesla, Meta)
  • $\square$
    I know the sender’s name (e.g., Elon Musk, Mark Zuckerberg)
  • $\square$
    The sender’s name and affiliated institution are consistent (e.g., Elon Musk & Tesla, Mark Zuckerberg & Meta)
  • $\square$
    I am acquainted with the sender (e.g., friend, student, professor)
Q4. [Recipient] Does the recipient name in the email align with your “real name”?
[…]
    Yes
[…]
    No
Q5. [Action] If you receive this email, what would be your action? (single select)
[…]
    Respond to this email
[…]
    Click the link
[…]
    Read the email and do nothing
[…]
    Move to spam
Q6. Do you think this email is genuine or phishing? (single select)
[…]
    Definitely genuine
[…]
    Probably genuine
[…]
    Unsure
[…]
    Probably phishing
[…]
    Definitely phishing
Q7. (Respond if you answered that the email seems like phishing in Q6) What aspect of the email contributed to your perception of it being phishing? (select all that apply)
  • $\square$
    Information on specific events, schedules, etc. (time, location, etc.)
  • $\square$
    Recipient information (recipient’s name, etc.)
  • $\square$
    Sender information (sender’s name, institution, email address, etc.)
  • $\square$
    Information related to expertise or interests
  • $\square$
    Purpose of the email

### v3

@@ 9 Conclusion
In this work, we examine the emerging threats posed by LLM agents, particularly their potential use of web-based tools in cyberattacks that exploit private data. Specifically, we focused on three types of cyberattacks: PII collection, impersonation post generation, and spear phishing email generation. Our experimental results demonstrate that attackers can successfully automate these attacks using LLM agents, with web-based tools enhancing the performance of these attacks. Furthermore, our results report easy bypass of LLM safeguards at low-costs, making utilization of LLM agents in cyberattacks a viable option. These findings expose a significant vulnerability in current safeguards and underscore the urgent need for security measures to prevent the misuse of LLM agents.
Acknowledgments
This work was partly supported by Institute for Information & communications Technology Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2023-00215700, Trustworthy Metaverse: blockchain-enabled convergence research, 50%), (RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST), 20%), and (RS-2024-00436680, Global Research Support Program in the Digital Field program, 20%). It was also supported by Artificial intelligence industrial convergence cluster development project (10%) funded by the Ministry of Science and ICT(MSIT, Korea)&Gwangju Metropolitan City.
@@ Acknowledgments
This paper discusses and examines the potential threat of LLM agents in cyberattacks exploiting personal data, which can raise ethical considerations. However, we minimize these concerns by employing the following measures:
Aggregation of PIIs. For annotation purposes, these data were provided to annotators in Excel file format. Once the annotation process was completed, the annotators returned the annotated files to the principal investigator and deleted their local copies. After acceptance of our paper, all files containing personal information were deleted.
User Study. According to the local IRB’s protocol, our research is excluded from the review for the following reasons. First, we used only publicly available information and did not handle any sensitive data. Additionally, the emails evaluated contained no harmful content, minimizing any potential impact on participants. In the recruitment process, we obtained informed consent from participants, detailing the purpose of the research, the expected duration, procedures involved, anticipated benefits, personal information protection, and consent for collecting and using their provided email addresses. We also informed participants that the LLM agents could search the internet for their email addresses to gather information and use it to generate emails. We asked them to evaluate emails through a Google survey after notification. Each participant was exposed only to emails generated from their own email address. All participant-related information was destroyed after the experiments, and participants were informed of this process. Additionally, they were assured that their responses would be excluded from the analysis if they withdrew consent, ensuring robust protection of their personal information.
Responsible Disclosure Process. We plan to share our findings to the LLM service vendors. Specifically, we will share detailed prompts and raise concerns regarding the safeguard vulnerabilities of web-based tools. Additionally, we will request the implementation of a rule ensuring that when an LLM utilizes web crawling tools, it checks and adheres to robots.txt directives. Also, we strongly recommend that the PII providers from whom we collected data create a robots.txt file to prevent web crawlers from accessing pages containing personal information.
@@
Our study explores the potential threat posed by LLM agents in cyberattacks. We acknowledge that disclosing the exact prompts used for these attacks could enable malicious actors to replicate them. To mitigate the risk of misuse, we have submitted our code for review with dummy prompts accessible at this link [^7] https://zenodo.org/doi/10.5281/zenodo.13691327. We will provide the full source code only upon request for legitimate research purposes.
Appendix A Details for Figure 1
In Figure 1, the ‘LLM agent’ denotes the highest capability setting for each task. For tasks involving PII collection and impersonation, the term refers to WebNav agents, whereas for spear-phishing email generation, it corresponds to WebSearch agents. The performance metrics vary according to each task and are denoted by the axes in the figure.
For the PII collection task, the metric represents the average number of PII items successfully collected from CS professors. In the case of the impersonation post generation task, it is measured by the percentage of posts judged as authentic in a binary (yes/no) authenticity test. For the spear-phishing email generation task, it indicates the click-through rate of links contained within emails designed for general purposes.
Appendix B PII collection - Human annotation
@@
Appendix C Impersonation Post Evaluation Templates
In Section 5, LLM evaluators were utilized in both the A/B Test and the Yes/No Test. The prompts provided are detailed below. These prompts were designed to enable LLMs to incorporate personal information, similar to how human evaluators use factual details like occupation and interests to assess post authenticity. This approach is particularly relevant in the context of advanced impersonation attacks, which exploit personal data to deceive effectively. Experimentally, we observed that without such prompts, LLMs tend to rely on fluency, favoring polished entries while failing to evaluate authenticity. By guiding models, we could align their evaluation process more closely with human judgment.
@@ Appendix D Approachability of Cyberattacks
Table 5 represents a cost analysis of PII collection, impersonation post generation, and phishing email generation using WebSearch agents.
@@ Appendix E Spear Phishing Email Generation
E.1 User Study
@@ E.2 Chi-Squared Tests
This section investigates the factors influencing participants’ perception of email authenticity. First, we identify which aspects of an email had the most significant influence on determining the email’s authenticity. We conducted chi-squared tests to evaluate the independence between email evaluation questions (Q1–Q4) and the question assessing perceived email authenticity (Q6). For details on the questions used in our survey, please refer to the document titled ‘Questionnaire for User Study,’ accessible at this link [^8] https://zenodo.org/doi/10.5281/zenodo.13691327. In Table 6, the results for “content detail” and “content authenticity” show p-values significantly less than 0.05, indicating strong evidence against the null hypothesis.
Further analysis focused on “content authenticity” (Q2), which was identified as the most strongly associated factor. Specifically, we examined which aspects of email content influenced participants’ perceptions of authenticity. We evaluated the independence between each response option within Q2 and the option in Q6 using chi-squared tests. The “Specific area of expertise” option had a p-value much less than 0.05, indicating a highly significant association.
@@ E.4 Contingency Table
Table 6: Chi-squared test results: Comparisons of Q1-Q4 and Q6, and response options within Q2 and Q6.
Questions Response Options within Q2
Q1 Q2 Q3 Q4 institution general expertise specific expertise hobby
(content detail) (content auth) (sender) (recipient)
kappa 98.80 113.54 56.54 3.69 4.34 0.01 11.15 0.28
p-value $2.23\times 10^{-5}$ $3.12\times 10^{-7}$ 0.0159 0.4497 0.0595 0.0373 0.9436 0.0008